OpenStudy (anonymous):

Okay, I'll try again. It was x-y=-3. (7th grade math)

Nnesha (nnesha):

solve for x or y ??? question ??

OpenStudy (anonymous):

The directions are " Find the x-intercept and y-intercept of each line " so i m confused on how to solve for x or y.

OpenStudy (kendricklamar2014):

Lets solve for X first @ihatespiders

OpenStudy (kendricklamar2014):

First, add y to both sides

OpenStudy (kendricklamar2014):

x - y + y = -3 + y

OpenStudy (kendricklamar2014):

Can you solve that ^ @ihatespiders

OpenStudy (anonymous):

i got x=-2

OpenStudy (kendricklamar2014):

Wrong, first solve that ^^

OpenStudy (kendricklamar2014):

x - y + y = leaves only x because -y + y cancels out the y's

OpenStudy (kendricklamar2014):

-3 + y = y - 3

OpenStudy (kendricklamar2014):

So for the X-intercept, X = Y - 3

OpenStudy (kendricklamar2014):

Do you get it so far @ihatespiders

OpenStudy (anonymous):

Hi - she got frustrated and walked away. Thanks for your help thought!

Nnesha (nnesha):

x-intercept when y = 0 so for intercept you have to replace y by 0 and then solve

Nnesha (nnesha):

y-intercept when x=0 so to find y -intercept replace x by 0 and then solve can you do it now ??

Nnesha (nnesha):

 for x-intercept now solve for x